# SemPLeS: Semantic Prompt Learning for Weakly-Supervised Semantic Segmentation

### v2

@@ 3.2 Semantic Prompt Learning for WSSS (SemPLeS)
Here, $\lambda$ is the weight for the refinement loss. It can be seen that, with the derived background prompts $p_{k}$ (fixed here) and the introduced refinement loss $L_{refine}$, the class-associated background regions would be suppressed from the foreground mask $M$, preventing possible false activation. More importantly, by jointly applying the matching and refinement objectives with image-level supervision, our SemPLeS framework advances vision-language learning to enhance the semantic alignment between the segmented regions and the target object categories, resulting in compact and complete object masks $M^{\prime}$ desired for segmentation. It is worth noting that, the CLIP model and the learned prompts $p_{k}$ are leveraged to guide the learning of the mask generator $S$ in our framework, and hence only the mask generator $S$ is needed for producing object masks $M^{\prime}$ in the WSSS pipeline when the training is complete.
@@ 4.2 Implementation Details
We adopt the Transformer-based backbone model in [82] for our mask generator $S$. For CLIP [50], we use ViT-B/32 [18] as the image encoder. The learnable prompts are randomly initialized with the sequence length $K=30$ and the prompt embedding dimension $D=512$. We adopt the cosine similarity where non-positive scores are clamped to a small positive number $0.0001$ following CLIMS [64]. For loss weights, we set $\lambda_{b}$, $\lambda_{T}$ and $\lambda$ as $2.4$, $0.02$ and $0.05$ for PASCAL VOC 2012 and $0.75$, $0.01$ and $0.2$ for MS COCO 2014. The default batch size is $64$. We set the initial learning rate to 5e-4 and 5e-6 and train our framework for $60$ epochs on PASCAL VOC 2012 and MS COCO 2014, respectively. The AdamW optimizer is adopted with the cosine scheduler. All models are implemented in PyTorch and trained on NVIDIA V100 GPUs.
4.3 Quantitative Comparisons
Table 1: Quantitative comparisons of CAMs (Seed) and the resulting pseudo masks (Mask) on PASCAL VOC 2012 train set.
Method Seed Mask
@@ 4.3 Quantitative Comparisons
$\text{MCTformer}_{\text{~{}~{}CVPR'22}}$ [67] 61.7 69.1
@@
$\text{WeakTr}_{\text{~{}~{}arXiv'23}}$ [82] 65.9 74.2
$\text{POLE}_{\text{~{}~{}WACV'24}}$ [48] 59.0 74.2
SemPLeS (Ours) 68.7 78.4
To evaluate our proposed SemPLeS framework, we first compare the quality of the initial activation maps (Seed) and also the resulting pseudo masks (Mask) with previous works, as shown in Table 1. From the results in this table, we see that our SemPLeS achieves the best performance compared with previous weakly-supervised segmentation methods. Specifically, our initial activation maps (Seed) achieve 68.7% and the produced pseudo masks (Mask) report 78.4% in mIoU. This verifies that, by exploiting CLIP to perform vision-language learning plus the designed prompt learning, our proposed SemPLeS framework successfully generates pixel-wise predictions from image-level supervision, which helps learn the following segmentation model.
For the final segmentation results, our SemPLeS achieves the best performance on PASCAL VOC and reports 83.4% and 82.9% mIoU on the validation and testing sets, respectively, as shown in Table 2. Our method outperforms the previous SOTA, FMA-WSSS [71], by $0.8\%$ and $1.3\%$ mIoU on the validation and testing sets, respectively. In addition, our SemPLeS achieves the SOTA performance of 56.1% mIoU on MS COCO. The above results verify that our method is effective in performing semantic segmentation from image-level supervision.
@@
CNN-/Transformer-based approaches
$\text{CLIMS}_{\text{~{}~{}CVPR'22}}$ [64] DL2-Res101 Res50 ${}^{\dagger}$ ${}^{\dagger}$
$\text{MCTformer}_{\text{~{}~{}CVPR'22}}$ [67] DL1-WRes38 DeiT-S 71.9 71.6 42.0
@@
$\text{OCR}_{\text{~{}~{}CVPR'23}}$ [16] DL1-WRes38 DeiT-S 72.7 72.0 42.5
$\text{MMCST}_{\text{~{}~{}CVPR'23}}$ [68] DL1-WRes38 ViT-B 72.2 72.2 45.9
$\text{LPCAM}_{\text{~{}~{}CVPR'23}}$ [12] DL1-WRes38 DeiT-S 72.6 72.4 42.8
$\text{FPR}_{\text{~{}~{}ICCV'23}}$ [8] DL2-Res101 Res50 70.3 70.1 43.9
$\text{USAGE}_{\text{~{}~{}ICCV'23}}$ [49] DL1-WRes38 DeiT-S 71.9 72.8 42.7
[…]
$\text{WeakTr}_{\text{~{}~{}arXiv'23}}$ [82] Sgmt-DeiT-S DeiT-S 73.2 74.0 44.4
$\text{MCC}_{\text{~{}~{}WACV'24}}$ [62] DeiT-B DeiT-B 70.3 71.2 42.3
$\text{POLE}_{\text{~{}~{}WACV'24}}$ [48] DL2-Res101 Res50 ${}^{\dagger}$ ${}^{\dagger}$
$\text{SFC}_{\text{~{}~{}AAAI'24}}$ [77] DL2-Res101 Res50 71.2 72.5 46.8
SAM-based approaches
$\text{SEPL}_{\text{~{}~{}NeurIPSW'23}}$ [10] DL2-Res101 Res50 71.1 - -
@@
[…]
SemPLeS (Ours) M2F-Swin-L ViT-B 83.4 82.9 56.1
@@
Comparison with CLIP-based and SAM-based methods: In light of the constraint in WSSS (i.e., training using only class labels), several approaches have leveraged CLIP to enhance the quality of the produced CAMs by prompting, such as CLIMS [64], CLIP-ES [41], MMCST [68], and POLE [48]. However, they either consider only the foreground class prompts, or rely on general background prompts defined by additional manual efforts and human knowledge. Moreover, such manually-defined prompts may not fully exploit the knowledge learned in CLIP. In contrast, our method automatically learns prompts embedded with class-associated semantic knowledge from the CLIP latent space with no need of any manual efforts, resulting in better performance than other CLIP-based methods, as shown in Table 2.
@@
Recently, SAM has been proposed to produce high-quality class-agnostic masks with overwhelming generalizability. Several approaches [10, 59, 13, 25] have explored the potential of leveraging SAM for WSSS, and most of them require additional foundation models (e.g., BLIP-2 [36], Grounding-DINO [43], and RAM [74]) to incorporate semantic information for semantic segmentation. Finally, FMA-WSSS [71] exploits both CLIP and SAM and achieves the SOTA performance among the above SAM-based methods, and we further outperform FMA-WSSS with the proposed SemPLeS framework, as shown in Table 2.
Compatibility with other WSSS methods: We evaluate the compatibility of our method by integrating the proposed SemPLeS framework with other WSSS methods, including CNN-based [64], Transformer-based [82], and SAM-based ones [71]. The quantitative results in Table 3 show that our method consistently improves all types of methods, demonstrating its compatibility and robustness. Moreover, our results show small margins to the fully-supervised upper bound (i.e., $2.6\%$ and $3.2\%$ with Mask2Former on the validation and testing set, respectively), indicating that the proposed SemPLeS can help diminish the performance gap between image-level and pixel-level supervision.
Table 3: Quantitative results of our proposed SemPLeS framework with different WSSS methods on PASCAL VOC 2012 val and test sets, including CNN-based (CLIMS [64]), Transformer-based (WeakTr [82]), and SAM-based (FMA-WSSS [71]) methods. “Seg. Model” and “Backbone” denote the segmentation network and the CAM generation network, respectively. “DL2”, “Sgmt” and “M2F” denote Deeplab-V2 [5], Segmenter [58], and Mask2Former [15], respectively. $\mathcal{FS}$ denotes full supervision, i.e., training the corresponding segmentation model with the ground truth masks.
Method Seg. Model Backbone val test
$\text{CLIMS}_{\text{~{}~{}CVPR'22}}$ [64] DL2-Res101 Res50 70.4 70.0
+SemPLeS (Ours) DL2-Res101 Res50 71.0 (+0.6) 71.5 (+1.5)
$\mathcal{FS}$ DL2-Res101 - 77.7 79.7
$\text{WeakTr}_{\text{~{}~{}arXiv'23}}$ [82] Sgmt-DeiT-S DeiT-S 73.2 74.0
+SemPLeS (Ours) Sgmt-DeiT-S DeiT-S 74.2 (+1.0) 74.8 (+0.8)
$\mathcal{FS}$ Sgmt-DeiT-S - 79.7 79.6
[…]
+SemPLeS (Ours) M2F-Swin-L ViT-B 83.4 (+0.8) 82.9 (+1.3)
$\mathcal{FS}$ M2F-Swin-L - 86.0 86.1
@@ 4.4 Qualitative Comparisons
Figure 3: Qualitative comparisons of the initial CAMs. “GT” denotes the ground truth masks. Note that our SemPLeS successfully suppresses the backgrounds and produces precise CAMs compared with WeakTr [82] and CLIMS [64] (which relies on manually-defined background prompts involved with the heuristic human knowledge).
Figure 4: Qualitative comparisons of segmentation maps with recent WSSS methods.
Figure 5: Visualization of the corresponding image regions of the manually-defined background prompts (determined in CLIMS [64]) and our learned prompts. For manual prompts, we select those with the highest similarity, lowest similarity, and also aggregated one (“All BG Prompts”) from the second to fourth columns, respectively.
For qualitative comparisons, our method shows more precise activation maps on various object categories and performs favorably compared with CLIMS [64] and WeakTr [82], as shown in Fig. 3. This validates that, by advancing image-text contrastive learning with learnable prompts, our SemPLeS would enhance the alignment between the segment regions and the target object categories. In Fig. 4, we also show that our final segmentation results are superior to recent WSSS methods. In addition, we also visualize the corresponding regions of our learnable prompts by calculating the similarities to image patches with the text and image encoders of CLIP, as shown in Fig. 5. The manually-defined background prompts [64] may falsely highlight the foreground objects due to their high co-occurrence when pre-training CLIP. Furthermore, such manual prompts are limited to specific categories, causing incorrect backgrounds to input images (e.g., the prompt “a photo of track" in the first example (cow)). In contrast, our learnable prompts successfully highlight all the background regions associated with each object category, demonstrating the effectiveness of our Contrastive Prompt Learning. It is also worth noting that, our class-associated prompts can also learn knowledge shared across object categories, such as the grass and tree region in both the first (cow) and third (dog) example in Fig. 5.
4.5 Ablation Studies
Table 4: Quantitative ablation studies of our loss functions on PASCAL VOC 2012.
@@ 4.5 Ablation Studies
Loss Component Analysis:
To analyze the importance of the introduced loss functions, we conduct both quantitative and qualitative ablation studies, as shown in Table 4 and Fig. 6. In Table 4, we see that applying only the matching loss $L_{match}$ would result in $67.6\%$ mIoU. If we directly add the refinement loss $L_{refine}$ without prompt learning, the performance would be similar since the prompts are randomly initialized and are not learned. When prompt learning is further considered, applying only the $L^{T}_{prompt}$ to repel the text labels may result in trivial solutions with little improvement. On the other hand, if only $L^{I}_{prompt}$ is enforced to align with the background images, the prompts are still likely to capture the semantics of the foreground object categories, resulting in $67.9\%$ mIoU. Finally, when $L^{I}_{prompt}$ and $L^{T}_{prompt}$ are jointly applied to learn the background regions while avoiding describing the foreground object categories, the mIoU would improve to $68.7\%$. Together with the qualitative results in Fig. 6, we validate that our proposed SemPLeS framework and semantic prompt learning would prevent false activation of co-occurring backgrounds and therefore benefit segmentation in a weakly-supervised fashion.
Figure 6: Qualitative ablation studies of loss functions. With $L_{refine}$ applied with our learned prompts, we see that our proposed SemPLeS framework successfully suppresses co-occurring backgrounds associated with distinct object categories.
Table 5: Loss weight analysis.
$\lambda$ 0 0.01 0.02 0.05 0.1
mIoU 67.6 68.3 68.4 68.7 67.8
Loss Weight Analysis:
In Table 5, we perform hyper-parameter analysis on the refinement loss weight $\lambda$ in Eq. 3. $\lambda=0$ means only $L_{match}$ is used. When varying $\lambda$, the mIoU gradually increases from $67.6\%$ to $68.7\%$ when $\lambda=0.05$, and then drops to $67.8\%$ when $\lambda=0.1$. Hence, we set $\lambda=0.05$ by default.
Table 6: Learning scheme variants. “seq.”: sequential, “alter.”: alternate.
Scheme # of rounds mIoU
joint training - 65.9
seq./alter. training 1 68.7
2 68.5
3 68.4
Learning Scheme Variants:
In Table 6, we provide quantitative analysis on PASCAL VOC 2012 when using different learning schemes in our proposed SemPLeS framework. We see that jointly training all the learning components in Fig. 2 would result in an inferior performance of $65.9\%$. This is because, in the beginning of training, the prompts are not well learned and therefore would not be desired for our (c) Prompt-guided Semantic Refinement. If we instead perform sequential training (i.e., $1$ round), the mIoU would improve to $68.7\%$. Since alternate training between (b) Contrastive Prompt Learning and (c) Prompt-guided Semantic Refinement in our SemPLeS framework would increase the training time and may lead to overfitting, we choose to sequentially train our SemPLeS framework for one round.
@@ 5 Conclusion
In this paper, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework, which advances vision-language learning to achieve weakly-supervised semantic segmentation (WSSS). In addition to exploiting the pre-trained CLIP model to perform Segment-Label Matching, we further present Contrastive Prompt Learning and Prompt-guided Semantic Refinement in the proposed SemPLeS framework to prevent false activation of image backgrounds. With no need to manually define background texts through prompt engineering, our learned prompts properly capture and suppress co-occurring backgrounds for each object category, resulting in precise activation maps for segmentation in a weakly-supervised fashion. Quantitative experiments on the segmentation benchmarks confirm the effectiveness of our proposed SemPLeS framework, and visualization and ablation studies are conducted to demonstrate and verify the interpretability of learned prompts. Our method achieves SOTA performance on the standard WSSS benchmarks, PASCAL VOC and MS COCO, and shows compatibility with other WSSS methods.

### v3

@@ 3.2.3 Prompt-guided Semantic Refinement
Here, $\lambda$ is the weight for the refinement loss. It can be seen that, with the derived background prompts $p_{k}$ (fixed here) and the introduced refinement loss $L_{refine}$, the class-associated background regions would be suppressed from the foreground mask $M$, preventing possible false activation. More importantly, by jointly applying the matching and refinement objectives with image-level supervision, our SemPLeS framework advances vision-language learning to enhance the semantic alignment between the segmented regions and the target object categories, resulting in compact and complete object masks $M^{\prime}$ desired for WSSS. It is worth noting that, the CLIP model and the learned prompts $p_{k}$ are leveraged to guide the learning of the mask generator $S$ in our framework, and hence only the mask generator $S$ is needed for producing object masks $M^{\prime}$ in the WSSS pipeline when the training is complete.
@@ 4.2 Implementation Details
For CLIP [51], we use ViT-B/32 [18] as the image encoder. Following [65], we adopt the cosine feature similarity where non-positive scores are clamped to a small positive number. The learnable prompts are randomly initialized with the sequence length $K=30$. The default batch size is $64$. We set the initial learning rate to 5e-4 and 5e-6 and train our framework for $60$ epochs on PASCAL VOC 2012 and MS COCO 2014, respectively. For loss weights, we set $\lambda_{b}$, $\lambda_{T}$ and $\lambda$ as $2.4$, $0.02$ and $0.05$ for PASCAL VOC 2012 and $0.75$, $0.01$ and $0.2$ for MS COCO 2014. The AdamW optimizer is adopted with the cosine scheduler. The proposed framework is implemented in PyTorch and trained with NVIDIA V100 GPUs.
Method CAM Mask
$\text{MCTformer}_{\text{~{}~{}CVPR'22}}$ [68] 61.7 69.1
@@
$\text{WeakTr}_{\text{~{}~{}arXiv'23}}$ [84] 65.9 74.2
@@
$\text{MCC}_{\text{~{}~{}WACV'24}}$ [62] - 73.0
$\text{POLE}_{\text{~{}~{}WACV'24}}$ [49] 59.0 74.2
$\text{DuPL}_{\text{~{}~{}CVPR'24}}$ [64] - 76.0
SemPLeS (Ours) 68.7 78.4
Table 1: Quantitative results of CAMs (CAM) and the resulting pseudo masks (Mask) on PASCAL VOC 2012 train set.
4.3 Quantitative Comparisons
Figure 2: Qualitative results of CAMs. “GT” denotes the ground truth masks. We see that our proposed SemPLeS framework produces precise CAMs better aligned with the ground truth masks.
Method Backbone val test
CNN-/Transformer-based approaches
SIPE CVPR’22 ${}_{\text{~{}~{}CVPR'22}}$ [9] DL2-Res101 68.8 69.7
$\text{CLIMS}_{\text{~{}~{}CVPR'22}}$ [65] DL2-Res101 70.4 70.0
$\text{MCTformer}_{\text{~{}~{}CVPR'22}}$ [68] DL1-WRes38 71.9 71.6
@@ 4.3 Quantitative Comparisons
$\text{MMCST}_{\text{~{}~{}CVPR'23}}$ [69] DL1-WRes38 72.2 72.2
$\text{LPCAM}_{\text{~{}~{}CVPR'23}}$ [12] DL1-WRes38 72.6 72.4
$\text{FPR}_{\text{~{}~{}ICCV'23}}$ [5] DL2-Res101 70.3 70.1
[…]
$\text{MCC}_{\text{~{}~{}WACV'24}}$ [62] DeiT-B 70.3 71.2
$\text{POLE}_{\text{~{}~{}WACV'24}}$ [49] DL2-Res101 71.5 71.4
$\text{SFC}_{\text{~{}~{}AAAI'24}}$ [79] DL2-Res101 71.2 72.5
$\text{DuPL}_{\text{~{}~{}CVPR'24}}$ [64] ViT-B 73.3 72.8
SAM-based approaches
$\text{SEPL}_{\text{~{}~{}NeurIPSW'23}}$ [10] DL2-Res101 71.1 -
@@
SemPLeS (Ours) DL2-Res101 73.9 73.8
SemPLeS (Ours) M2F-Swin-L 83.4 82.9
Table 2: Quantitative results of segmentation masks on PASCAL VOC 2012 [20] val and test sets. “Backbone” denotes the segmentation network. “DL”, “Res”, “WRes”, and “M2F” denote DeepLab [6], ResNet [23], WideResNet [74], and Mask2Former [15], respectively.
Method COCO val
$\text{MCC}_{\text{~{}~{}WACV'24}}$ [62] 42.3
$\text{USAGE}_{\text{~{}~{}ICCV'23}}$ [50] 42.7
$\text{LPCAM}_{\text{~{}~{}CVPR'23}}$ [12] 42.8
$\text{FPR}_{\text{~{}~{}ICCV'23}}$ [5] 43.9
@@
$\text{WeakTr}_{\text{~{}~{}arXiv'23}}$ [84] 44.4
$\text{DuPL}_{\text{~{}~{}CVPR'24}}$ [64] 44.6
@@
SemPLeS (Ours) 56.1
Table 3: Quantitative results of the segmentation masks on MS COCO 2014 [41] val set.
To evaluate our proposed SemPLeS framework, we follow the standard WSSS pipeline and take our refined masks $M^{\prime}$ as CAMs to produce pseudo masks. In Table 1, we compare the quality of CAMs and also the resulting pseudo masks with previous works. From the results in this table, we see that our SemPLeS achieves the best performance compared with previous weakly-supervised segmentation methods. Specifically, our CAMs achieve 68.7% and the produced pseudo masks report 78.4% in mIoU. This verifies that, by exploiting CLIP to perform vision-language learning plus the designed prompt learning, our proposed SemPLeS framework successfully generates pixel-wise predictions from image-level supervision, which helps learn the following segmentation network.
In Table 2, by taking the derived pseudo masks to train the segmentation networks, we see that our SemPLeS achieves the best performance on PASCAL VOC and reports $83.4\%$ and $82.9\%$ mIoU on the validation and testing sets, respectively. Our method outperforms the previous work, FMA-WSSS [72], by $0.8\%$ and $1.3\%$ mIoU on the validation and testing sets, respectively. In addition, our SemPLeS achieves the competitive performance of $56.1\%$ mIoU on MS COCO in Table 3. The above results verify that our method is effective in performing semantic segmentation from image-level supervision.
Figure 3: Qualitative results of segmentation maps. “GT” denotes the ground truth masks.
Method val test
CLIMS [65] 70.4 70.0
+SemPLeS (Ours) 71.0 71.5
WeakTr [84] 73.2 74.0
+SemPLeS (Ours) 74.2 74.8
FMA-WSSS [72] 82.6 81.6
+SemPLeS (Ours) 83.4 82.9
Table 4: Quantitative results of our proposed SemPLeS framework based on different WSSS methods, including CNN- (CLIMS), Transformer- (WeakTr), and SAM-based (FMA-WSSS) ones.
Comparison with CLIP-based methods:
@@
In light of the constraint in WSSS (i.e., training using only class labels), several approaches have leveraged CLIP to enhance the quality of the produced CAMs by prompting, such as CLIMS [65], CLIP-ES [42], MMCST [69], and POLE [49]. However, they either consider only the foreground class prompts, or rely on general background prompts defined by additional manual efforts and human knowledge. Moreover, such manually-defined prompts may not fully exploit the knowledge learned in CLIP. In contrast, our method automatically learns prompts embedded with class-associated semantic knowledge from the CLIP latent space with no need of any manual efforts, resulting in better performance than these CLIP-based methods in Table 2.
Comparison with SAM-based methods:
@@
Recently, SAM has been proposed to produce high-quality class-agnostic masks with overwhelming generalizability. Several approaches [10, 59, 13, 26] have explored the potential of leveraging SAM for WSSS, and most of them require additional foundation models (e.g., BLIP-2 [37], Grounding-DINO [44], and RAM [76]) to incorporate semantic information for semantic segmentation. Specifically, FMA-WSSS [72] exploits CLIP and achieves competitive performance among the above methods, and we further outperform FMA-WSSS with the proposed SemPLeS framework, as shown in Table 2.
Compatibility with other WSSS methods:
We evaluate the compatibility of our method by integrating the proposed SemPLeS framework with other WSSS methods, including CNN-based [65], Transformer-based [84], and SAM-based ones [72]. The quantitative results are presented in Table 4. From this table, we see that our SemPLeS improves all types of methods, demonstrating the compatibility and robustness of the proposed framework. It is worth noting that, even though CLIMS [65] already uses manually-defined prompts, our SemPLeS could still achieve further improvement with our learnable prompts. This shows that manually-defined prompts are limited and may not fully exploit the CLIP latent space, while our learnable prompts is able to automatically capture the semantic knowledge associated with target object categories to complement such pre-defined prompts, verifying the effectiveness of our designed prompt learning and proposed SemPLeS framework.
@@ 4.4 Qualitative Comparisons
For qualitative comparisons, our method shows more precise activation maps on various object categories and performs favorably compared with previous works, as shown in Fig. 2. In Fig. 3, we also show that our segmentation results are superior to other WSSS methods. This validates that, by advancing image-text contrastive learning with learnable prompts, our SemPLeS would enhance the alignment between the segment regions and the target object categories, resulting in precise CAMs and segmentation masks better aligned with the ground truth.
In addition, we also visualize the corresponding regions of our learned prompts by calculating the similarities to image patches with the text and image encoders of CLIP. As shown in Fig. 4, the manually-defined background prompts [65] may falsely highlight the foreground objects (e.g., the bird example in the third row) due to their high co-occurrence when pre-training CLIP. Also, such manual prompts are limited and may fail to cover the whole background in images (e.g., the cow example in the first row). In contrast, our learned prompts highlight all the background regions associated with each object category, showing the effectiveness of our Contrastive Prompt Learning.
Figure 4: Visualization of the manually-defined background prompts [65] and our learned prompts.
@@
Table 5: Quantitative ablation studies of our loss functions. With both $L^{T}_{prompt}$ and $L^{I}_{prompt}$ applied (Eq. (2)), the derived prompts would be desired to guide the semantic refinement through loss $L_{refine}$, resulting in the best performance.
4.5 Ablation Studies
To analyze the importance of the introduced loss functions, we conduct both quantitative and qualitative ablation studies, as shown in Table 5 and Fig. 5. In Table 5, we see that applying only the matching loss $L_{match}$ would result in $67.6\%$ mIoU. If we directly add the refinement loss $L_{refine}$ without prompt learning, the performance would be similar since the prompts are randomly initialized and are not learned. When prompt learning is further considered, applying only the $L^{T}_{prompt}$ to repel the text labels may result in trivial solutions with little improvement. On the other hand, if only $L^{I}_{prompt}$ is enforced to align with the background images, the prompts are still likely to capture the semantics of the foreground object categories, resulting in $67.9\%$ mIoU. Finally, when $L^{I}_{prompt}$ and $L^{T}_{prompt}$ are jointly applied to learn the background regions while avoiding describing the foreground object categories, the mIoU would improve to $68.7\%$. Together with the qualitative results in Fig. 5, we validate that our designed prompt learning and the proposed SemPLeS framework would prevent false activation of co-occurring backgrounds and therefore benefit segmentation in a weakly-supervised fashion.
Figure 5: Qualitative ablation studies of loss functions. With both $L_{prompt}$ and $L_{refine}$ applied, the resulting CAMs are better aligned with the ground truth masks.
@@ 5 Conclusion
In this paper, we propose a Semantic Prompt Learning for WSSS (SemPLeS) framework, which advances vision-language learning to achieve weakly-supervised semantic segmentation (WSSS). In addition to exploiting the pre-trained CLIP model to perform Segment-Label Matching, we further present Contrastive Prompt Learning and Prompt-guided Semantic Refinement in the proposed SemPLeS framework to prevent false activation of image backgrounds. With no need to manually define background texts through prompt engineering, our learned prompts properly capture and suppress co-occurring backgrounds for each object category, resulting in precise activation maps for segmentation in a weakly-supervised fashion. Quantitative experiments on the segmentation benchmarks confirm the effectiveness of our proposed SemPLeS framework, and visualization and ablation studies are conducted to demonstrate and verify the effectiveness of learned prompts. Our method achieves competitive performance on the standard WSSS benchmarks, PASCAL VOC 2012 and MS COCO 2014, and shows compatibility with other WSSS methods.
6 Appendix
Figure 6: Failure case analysis. From the first to the third row, we show failure cases where the objects are partially visible, of small size, and visually similar to the surroundings, respectively.
6.1 Failure Case Analysis
In Fig. 6, we show several types of failure cases. In the first row, the potted plant in the top right corner is only partially visible in the image and thus is not easily recognized. In the second row, the two people in the top left corner are far away from the camera. The reduced size of such distant objects result in the failure of segmentation. As for the third row, the potted plant on the left is visually similar to its surroundings and therefore may confuse the models. Note that our SemPLeS still outperforms the previous SOTA (FMA-WSSS) in these challenging cases.
6.2 Limitations and Potential Negative Impact
Limitations:
In weakly-supervised semantic segmentation (WSSS), as only image-level labels are available for training, existing WSSS methods still struggle to segment the objects that are partially visible, of small size, or visually similar to the surroundings, as shown in Fig. 6.
Potential negative impact:
No specific potential negative impact is identified in this work.